# thanks: Authors contributed equallythanks: Authors contributed equallythanks: Authors contributed equally

### v1

@@ Abstract
Quantum error correction (QEC) provides a practical path to fault-tolerant quantum computing through scaling to large qubit numbers, assuming that physical errors are sufficiently uncorrelated in time and space. In superconducting qubit arrays, high-energy impact events produce correlated errors, violating this key assumption. Following such an event, phonons with energy above the superconducting gap propagate throughout the device substrate, which in turn generate a temporary surge in quasiparticle (QP) density throughout the array. When these QPs tunnel across the qubits’ Josephson junctions, they induce correlated errors. Engineering different superconducting gaps across the qubit’s Josephson junctions provides a method to resist this form of QP tunneling. By fabricating all-aluminum transmon qubits with both strong and weak gap engineering on the same substrate, we observe starkly different responses during high-energy impact events. Strongly gap engineered qubits do not show any degradation in $T_{1}$ during impact events, while weakly gap engineered qubits show events of correlated degradation in $T_{1}$. We also show that strongly gap engineered qubits are robust to QP poisoning from increasing optical illumination intensity, whereas weakly gap engineered qubits display rapid degradation in coherence. Based on these results, gap engineering removes the threat of high-energy impacts to QEC in superconducting qubit arrays.
Scalable and useful quantum computing is achievable through the use of quantum error correction (QEC), provided the hardware satisfies key underlying assumptions. One such assumption is the absence of any substantial source of correlated error. In superconducting qubit arrays, this assumption is heavily violated by error bursts due to high-energy impacts [1, 2, 3, 4, 5, 6, 7]. During an impact event, high-energy radiation deposits up to 1 MeV in the device substrate, which radiates throughout the device as high-energy phonons [1, 8]. As illustrated in Fig. 1(a), these phonons then break Cooper pairs, producing significantly elevated quasiparticle (QP) densities. These QPs tunnel through the Josephson junction, absorbing the qubit excitation and inducing excess $T_{1}$ decay [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Impacts therefore produce a massive correlated error burst that extends through time and space, which is not correctable using standard QEC. The presence of these bursts induces an error floor on the logical performance of QEC experiments, preventing the promised exponential suppression of logical error with increasing code size [20, 21]. Without mitigation in hardware, this effect single-handedly prevents scaling QEC to algorithmically relevant logical error rates.
@@
Meanwhile, gap engineering has been shown to modify the dynamics of QP tunneling and improve the charge-parity lifetimes in superconducting qubits [9, 10, 27, 28, 29, 30, 31, 32, 33, 34, 35]. It has also been shown to improve the speed of recovery from QP poisoning from high-energy impact events [6]. Our approach to gap engineering involves producing different superconducting gaps on each side of the junction presents an energy barrier to QPs, as illustrated in Fig. 1(b). The superconducting gap in thin films of aluminum is sensitive to thickness, with thinner films displaying higher superconducting gaps. When the gap difference between the thin and thick leads of the Josephson junction ($\delta\Delta=\Delta_{\mathrm{thin}}-\Delta_{\mathrm{thick}}$) is significantly larger than the qubit energy ($E_{q}=hf_{q}$), QPs with energy near the gap can no longer absorb the qubit energy required to climb up the barrier to the higher gap material. This suppresses the primary mechanism by which QPs induce $T_{1}$ decay.
@@ I.1 Gap engineering at the Junction
The thin junction lead provides a high-gap region separating the junction from the low-gap material in the qubit capacitor. In a strongly gap engineered qubit, this barrier exceeds the energy scale of the qubit. QPs near the superconducting gap in either capacitor pad then have insufficient energy to overcome the barrier and therefore cannot tunnel across the junction. As such, strongly gap engineered qubits will suppress the QP tunneling process that produces severe qubit decay, removing the dominant impact of elevated QP density on qubit $T_{1}$. QPs should be mainly generated in the capacitor pads, and QPs in the thin layer should preferentially fall into the thick capacitor material, rapidly depleting any stray QPs in the thin layer. This should prevent any significant density reaching the junction lead, so we do not expect to need additional depositions to achieve strong suppression of QP tunneling.
@@ I.2 Device Details
We fabricate an array of 12 transmon qubits in an all-aluminum device. We form a checkerboard placement of weakly and strongly gap engineered qubits to ensure that each class of qubit is uniformly distributed across the device, as illustrated in Fig. 1(d). The qubits are frequency-tunable with maximum frequencies in the range $f_{\text{max}}\approx 7-7.3$ GHz. We optimize the qubit geometry for coherence in this test device following Refs.  [36, 14]. Each qubit footprint is approximately 0.6 mm $\times$ 0.8 mm, with junction leads that are 2 $\mu$m long and 150 nm wide. We provide further information about the device fabrication and circuit parameters in the Supplementary Materials.
We achieve different levels of gap engineering by choosing the thicknesses of the aluminum layers at the junction. In all qubits, we choose the thick layer to be 100 nm to approach the bulk gap value $\Delta_{\mathrm{bulk}}$ and allow us to select over a wide range of gap differences $\delta\Delta$. For “weakly gap engineered” qubits, we choose 30 nm for the thin layer. This thickness targets a gap difference of around 5 GHz, substantially lower than the qubit energy for standard operating parameters. This gap difference also allows us to flux-tune qubits to explore the regimes with qubit frequencies both above and below resonance with the gap difference. For “strongly gap engineered” qubits, we choose 15 nm for the thin layer, which targets a gap difference of around 12 GHz, well above the maximum qubit frequency.
@@ II Rapid Repetitive Correlated Sampling
We perform rapid repetitive correlated sampling as in Ref. [3]. This involves initializing all qubits in the first excited state |$\ket{1}$, idling for 1 $\mu$s and then measuring all qubits. If the qubit is measured to be in the ground state |$\ket{0}$, we consider that a qubit error. This cycle is repeated every 100 $\mu$s over the span of 60 seconds to form a single dataset of $600,000$ samples. Over 100 datasets taken sequentially (6000 seconds total), we find 154 high-energy impact events, corresponding to an average time between events of close to 40 seconds. We include further analysis in the Supplementary Material. The corresponding event flux ($2.6\mathrm{s}^{-1}\mathrm{cm}^{-2}$) matches well to the event flux found on larger devices ($2.1\mathrm{s}^{-1}\mathrm{cm}^{-2}$ [3]). As in Ref. [3], when we perform the experiment initializing |$\ket{0}$ considering we do not find any evidence of excess rate of qubits being excited to |$\ket{1}$ during impact events.
Fig. 2 shows a single dataset featuring six events. As in Ref. [3], we use a matched filter to locate events with high confidence. Events are clearly visible on the weakly gap engineered qubits. Correspondingly, the performance of these qubits deviates noticeably from the prediction from independent $T_{1}$ and readout errors. For instance, all six weakly gap engineered qubits experience an error simultaneously two orders of magnitude more frequently than expected, indicative of highly correlated error bursts. The same events do not affect the performance of the strongly gap engineered qubits, as they continue to match the independent error model. In all 100 datasets, we find no time period where the behaviour of the strongly gap engineered qubits deviate from the independent error model, including where we see clear evidence of impacts in the weakly gap engineered qubits.
Fig. 3 shows the time series data around a single impact event. This is the event labeled $4$ shown in Fig. 2. In Fig. 3(a), we see that the event affects only the weakly gap engineered qubits, elevating their error rate significantly. By contrast, the strongly gap engineered qubits do not respond to the event, displaying error rates expected from independent $T_{1}$ and readout fidelities. During the event, we see substantial variation in error, revealing the checkerboard pattern of strongly and weakly gap engineered qubits directly in Fig. 3(b).
@@ III Illumination of the sample
Fig. 4(b) shows $T_{1}$ on all the qubits at their idle frequencies, which are in the range of 6.3 GHz to 6.6 GHz, in response to varying magnitudes of optical power. As we increase the normalized optical power from 00 to $10^{-3}$, the weakly gap engineered qubits (orange circles) suffer from large $T_{1}$ reductions of over an order of magnitude. On the other hand, for the strongly gap engineered qubits (green circles), we find insignificant increase of the relaxation rates from QPs until the normalized optical power reaches $10^{-1}$. We find similar robustness in CPMG dephasing lifetimes $T_{2,\mathrm{CPMG}}$ for the strongly gap engineered qubits. The rapid reduction in qubit $T_{1}$ in the weakly gap engineered qubits limits our ability to measure $T_{2,\mathrm{CPMG}}$ during optical illumination for those qubits.

### v2

@@ Abstract
Quantum error correction (QEC) provides a practical path to fault-tolerant quantum computing through scaling to large qubit numbers, assuming that physical errors are sufficiently uncorrelated in time and space. In superconducting qubit arrays, high-energy impact events can produce correlated errors, violating this key assumption. Following such an event, phonons with energy above the superconducting gap propagate throughout the device substrate, which in turn generate a temporary surge in quasiparticle (QP) density throughout the array. When these QPs tunnel across the qubits’ Josephson junctions, they induce correlated errors. Engineering different superconducting gaps across the qubit’s Josephson junctions provides a method to resist this form of QP tunneling. By fabricating all-aluminum transmon qubits with both strong and weak gap engineering on the same substrate, we observe starkly different responses during high-energy impact events. Strongly gap engineered qubits do not show any degradation in $T_{1}$ during impact events, while weakly gap engineered qubits show events of correlated degradation in $T_{1}$. We also show that strongly gap engineered qubits are robust to QP poisoning from increasing optical illumination intensity, whereas weakly gap engineered qubits display rapid degradation in coherence. Based on these results, gap engineering mitigates the threat of high-energy impacts to QEC in superconducting qubit arrays.
Scalable and useful quantum computing is achievable through the use of quantum error correction (QEC), provided the hardware satisfies key underlying assumptions. One such assumption is the absence of any substantial source of correlated error. In superconducting qubit arrays, this assumption is heavily violated by error bursts due to high-energy impacts [1, 2, 3, 4, 5, 6, 7]. During an impact event, high-energy radiation deposits up to 1 MeV in the device substrate, which radiates throughout the device as high-energy phonons [1, 8]. As illustrated in Fig. 1(a), these phonons then break Cooper pairs, producing significantly elevated quasiparticle (QP) densities. When these QPs tunnel through the Josephson junction, they can absorb the qubit excitation and induce excess $T_{1}$ decay [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Impacts can therefore produce a massive correlated error burst that extends through time and space, which is not correctable using standard QEC. The presence of these bursts induces an error floor on the logical performance of QEC experiments, preventing the promised exponential suppression of logical error with increasing code size [20, 21]. Without mitigation in hardware, this effect single-handedly prevents scaling QEC to algorithmically relevant logical error rates.
@@
Meanwhile, gap engineering has been shown to modify the dynamics of QP tunneling and improve the charge-parity lifetimes in superconducting qubits [9, 10, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. It has also been shown to improve the speed of recovery from QP poisoning from high-energy impact events [6]. Our approach to gap engineering involves producing different superconducting gaps on each side of the junction presents an energy barrier to QPs, as illustrated in Fig. 1(b). The superconducting gap in thin films of aluminum is sensitive to thickness, with thinner films displaying higher superconducting gaps. When the gap difference between the thin and thick leads of the Josephson junction ($\delta\Delta=\Delta_{\mathrm{thin}}-\Delta_{\mathrm{thick}}$) is significantly larger than the qubit energy ($E_{q}=hf_{q}$), QPs with energy near the gap can no longer absorb the qubit energy required to climb up the barrier to the higher gap material. This suppresses the primary mechanism by which QPs induce $T_{1}$ decay.
@@ I.2 Device Details
We fabricate an array of 12 transmon qubits in an all-aluminum device. We form a checkerboard placement of weakly and strongly gap engineered qubits to ensure that each class of qubit is uniformly distributed across the device, as illustrated in Fig. 1(d). The qubits are frequency-tunable with maximum frequencies in the range $f_{\text{max}}\approx 7-7.3$ GHz and feature shared Purcell filters[37, 38]. We optimize the qubit geometry for coherence in this test device following Refs.  [39, 14]. Each qubit footprint is approximately 0.6 mm $\times$ 0.8 mm, with junction leads that are 2 $\mu$m long and 150 nm wide. We provide further information about the device fabrication and circuit parameters in the Supplementary Materials.
We achieve different levels of gap engineering by choosing the thicknesses of the aluminum layers. Each qubit consists of two layers of aluminum with different thicknesses, which overlap in the capacitor pads and at the Josephson junction. The junction leads consist of only one layer of aluminum, one lead being thick and the other thin. In all qubits, we choose the thick layer to be 100 nm to approach the bulk gap value $\Delta_{\mathrm{bulk}}$ and allow us to select over a wide range of gap differences $\delta\Delta$. For “weakly gap engineered” qubits, we choose 30 nm for the thin layer. This thickness targets a gap difference of around 5 GHz, substantially lower than the qubit energy for standard operating parameters. This gap difference also allows us to flux-tune qubits to explore the regimes with qubit frequencies both above and below resonance with the gap difference. For “strongly gap engineered” qubits, we choose 15 nm for the thin layer, which targets a gap difference of around 12 GHz, well above the maximum qubit frequency.
I.3 Gap-engineering at the junction
The thin junction lead provides a high-gap region separating the junction from the two regions of low-gap material - the thick junction lead on one side and the thick layer of the qubit capacitor pad on the other. In a strongly gap engineered qubit, this barrier exceeds the energy scale of the qubit. QPs near the superconducting gap in either capacitor pad then have insufficient energy to overcome the barrier and therefore either cannot reach or cannot tunnel across the junction. As such, strongly gap engineered qubits will suppress the QP tunneling process that produces severe qubit decay, removing the dominant effect of elevated QP density on qubit $T_{1}$. QPs should be mainly generated in the capacitor pads. QPs in the thin layer of the capacitor pads will fall into the lower-gap thick layer, rapidly depleting any stray QPs in the thin layer. This should prevent any significant density of QPs moving from the thin layer of the qubit capacitor pad into the thin junction lead, so we do not expect to need additional depositions to achieve strong suppression of QP tunneling.
We note that the simplicity of our approach relies on the fact that the thin aluminum layer is the highest-gap material in our qubits. If the capacitor pad featured higher-gap material than the aluminum thin junction lead (e.g. Nb or Ta capacitor pads), the barrier would be compromised. QPs in the capacitor pad on the thin lead side would always be able to flow into the lead and across the junction. In this case, the barrier could instead be made using a yet-higher-gap material.
@@ II Rapid Repetitive Correlated Sampling
We perform rapid repetitive correlated sampling as in Ref. [3]. This involves initializing all qubits in the first excited state |$\ket{1}$, idling for 1 $\mu$s and then measuring all qubits. If the qubit is measured to be in the ground state |$\ket{0}$, we consider that a qubit error. This cycle is repeated every 100 $\mu$s over the span of 60 seconds to form a single experiment of $600,000$ samples. The full dataset was produced by first independently measuring the T1 and readout fidelity on each qubit to inform the independent error model, followed by 100 correlated sampling experiments as described above. This procedure took around 8 hours total time and represents 6000 seconds total sampling time. As the typical timescale for variation in individual qubit T1 due to TLSs is only several hours [40], parameter drift can lead to small discrepancies between the independent model performance and the performance of qubits during the experiment. We find 154 high-energy impact events, corresponding to an average time between events of close to 40 seconds. We include further analysis in the Supplementary Material. The corresponding event flux ($2.6\mathrm{s}^{-1}\mathrm{cm}^{-2}$) matches well to the event flux found on larger devices ($2.1\mathrm{s}^{-1}\mathrm{cm}^{-2}$ [3]). As in Ref. [3], when we perform the experiment initializing |$\ket{0}$ considering we do not find any evidence of an excess rate of qubits being excited to |$\ket{1}$ during impact events.
Fig. 2 shows a data from a single 60 second experiment featuring six events. As in Ref. [3], we use a matched filter to locate events with high confidence. The matched filter exploits the reliable shape of the impact event in the error time-series, and is mathematically equivalent to a scaled cross-correlation with a template of the event shape. Events are clearly visible on the weakly gap engineered qubits. Correspondingly, the performance of these qubits deviates noticeably from the prediction from independent $T_{1}$ and readout errors. For instance, all six weakly gap engineered qubits experience an error simultaneously two orders of magnitude more frequently than expected, indicative of highly correlated error bursts. The same events do not affect the performance of the strongly gap engineered qubits, as they continue to match the independent error model. In all 100 datasets, we find no time period where the behaviour of the strongly gap engineered qubits deviate from the independent error model, including where we see clear evidence of impacts in the weakly gap engineered qubits.
Fig. 3 shows the time series data around a single impact event. This is the event labeled $4$ shown in Fig. 2. In Fig. 3(a), we see that the event affects only the weakly gap engineered qubits, elevating their error rate significantly. By contrast, the strongly gap engineered qubits do not respond to the event, displaying error rates expected from independent $T_{1}$ and readout fidelities. During the event, we see substantial variation in error, revealing the checkerboard pattern of strongly and weakly gap engineered qubits directly in Fig. 3(b). During all discovered impact events, we see the SGE qubit display no additional error compared to their performance away from events and matching the independent error model. This confirms that the SGE qubits are robust to QP-induced correlated errors from impact events.
@@ III Illumination of the sample
Fig. 4(b) shows $T_{1}$ on all the qubits at their idle frequencies, which are in the range of 6.3 GHz to 6.6 GHz, in response to varying magnitudes of optical power. As we increase the normalized optical power from 00 to $10^{-3}$, the weakly gap engineered qubits (orange circles) suffer from large $T_{1}$ reductions of over an order of magnitude. On the other hand, for the strongly gap engineered qubits (green circles), we find insignificant increase of the relaxation rates from QPs until the normalized optical power reaches $10^{-1}$. We find similar robustness in CPMG dephasing lifetimes $T_{2,\mathrm{CPMG}}$ for the strongly gap engineered qubits. The $T_{2}$ in the weakly gap engineered qubits is limited by the rapid reduction in qubit T1.